# HyperMono: A Monotonicity-aware Approach to Hyper-Relational Knowledge Representation

### v1

@@ 5.1.2 Baselines, Evaluation Protocol and Implementation Details
  • m-TransH [13] builds on TransH by transforming each hyper-relational fact using a star-to-clique conversion.
  • RAE [14] integrates m-TransH with a multi-layer perceptron by considering the relatedness of entities. However, since the distance-based scoring function of TransH enforces constraints on relations, it fails to represent some binary relations in KGs.
  • NaLP-Fix [15] uses a convolutional-based framework to compute a relatedness vector for a triple. However, the model does not distinguish between the main triple and relation-specific qualifiers.
  • HINGE [15] adopts a convolutional framework for modeling hyper-relational facts. However, it only operates on a triple-quintuple level that lacks the necessary granularity for representing a relation instance with its qualifiers.
  • ShrinkE [16] models the primary triplets as a spatial-functional transformation from the head into a relation-specific box.
@@
  • StarE [17] employs a message passing based graph encoder to obtain entity/relation embedings, and feeds these embedings to a transformer decoder to score hyper-relational facts.
  • QUAD [18] introduces a graph encoder that aggregates information from the perspective of the qualifier entities. There are two variants of QUAD: QUAD and QUAD (Parallel). If there is no special suffix, QUAD denotes QUAD (Parallel).
@@
  • HyTrans [19] replaces the graph encoder with light-weight embedding modules, achieving higher efficiency without sacrificing effectiveness.
@@
  • GRAN [20] represents a fact as a heterogeneous graph, and employs edge-specific attentive bias to capture both local and global dependencies within the given fact. Note that GRAN contains three variants, i.e., GRAN-hete is the full model, GRAN-homo retains the graph structure but ignores heterogeneity and GRAN-complete considers neither the graph structure nor heterogeneity. If there is no special suffix, GRAN denotes GRAN-hete.
@@
  • HyNT [21] proposes a unified framework that learns representations of a hyper-relational knowledge graph containing numeric literals in either triples or qualifiers.
@@
  • HAHE [22] presents a model with hierarchical attention in global and local level to encode the graphical structure and sequential structure via self-attention layers.
@@
  • HyperFormer [23] considers local-level sequential information to encode the content of entities, relations and qualifiers. It includes an entity neighbor aggregator, relation qualifier aggregator and convolution-based bidirectional interaction modules.
Evaluation Protocol. Following existing work [21, 16, 22, 23], we rank all entities to predict the head entity in query (?$,r,t,\mathcal{Q})$ or the tail entity in query $(h,r,\texttt{?},\mathcal{Q})$. Specifically, we employ MRR and Hits@$k$ (abbreviated sometimes as H@k, $k\in\{1,3,10\}$) as evaluation metrics. MRR defines the inverse of the rank for the first correct answer, while Hits@k calculates the percentage of correct types ranked among the top-k answers. For all metrics, the larger the value, the better. Following the standard evaluation protocol in the filtered setting [1], all true triples in the HKGs are filtered out during evaluation, since predicting a low rank for these triples should not be penalized. All metrics are computed by averageing over two directions: head entity prediction and tail entity prediction.
Implementation Details. We conduct all experiments on six 32G Tesla V100 GPUs, and employ the AdamW [44] optimizer and a cosine decay scheduler with a linear warm-up for optimization. The hyperparameters are tuned according to the MRR on the validation set. We use grid search to select the optimal hyperparameters, mainly including: the learning rate $lr\in$3e-4, 4e-4, 5e-4, 6e-4, 7e-4}, the label smooth values $ls\in$0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9}, the embedding dimension $d\in$50, 100, 150, 200, 400, 600}, the temperature of Equation (1) $\gamma\in$1, 2, 3, 4, 5, 6}, the initialization range of embeddings $ir\in$0.01, 0.02, 0.05, 0.1, 0.2, 0.35, 0.5}, the neighbors of the head entity {2, 3, 4}, the number of qualifiers for a relation {5, 6, 7}, the number of layers in the transformer {2, 4, 8}, the number of heads in transformer {1, 2, 4, 8}, the input dropout rate in the transformer {0.6, 0.7, 0.8}.
5.2 Main Results
To address RQ1, we conduct experiments on the three types of datasets described above. The corresponding results are shown in Tables II, III, and IV.
Mixed-percentage Mixed-qualifier. Table II summarizes the results on WD50K, WikiPeople, and JF17K under the mixed-percentage mixed-qualifier condition. We can observe that HyperMono outperforms existing SoTA baselines by a large margin across most metrics on WD50K and JF17K. In particular, when compared with HyperFormer (the best performing baseline), HyperMono respectively achieves 2.1% and 2.9% improvements in MRR and Hits@1 on the JF17K dataset. A similar phenomenon occurred on the WD50K dataset. However, for the WikiPeople dataset, even though we can achieve optimal performance on the Hits@10 metric, we cannot surpass the best performing baseline on MRR and Hits@1. The intuitive reason for this is that, in comparison to the WD50K and JF17K datasets where 13.6% and 45.9% of main triple respectively contain hyper-relational knowledge, only 2.6% of triples in WikiPeople have qualifier pairs. This hinders HyperMono’s ability to exploit stage reasoning and qualifier monotonicity. Specifically, we employ a scaling transformation in the cone space to capture qualifier monotonicity, but the lack of qualifier pairs constrains this functionality.
TABLE III: Evaluation of different models under the fixed-percentage mixed-qualifier scenario on the WD50K, WikiPeople and JF17K datasets. Best scores are highlighted in bold, the second best scores are underlined. QUAD denotes the QUAD (Parallel) model in Table II, and GRAN denotes the GRAN-hete model in Table II.
Methods WD50K WikiPeople JF17K
33% 66% 100% 33% 66% 100% 33% 66% 100%
MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1 MRR H@1
Embedding-based Methods
ShrinkE 0.246 0.190 0.356 0.294 0.470 0.404 0.176 0.129 0.244 0.198 0.332 0.276 0.259 0.176 0.263 0.185 0.299 0.220
Transformer-based with GNNs Methods
StarE 0.308 0.247 0.449 0.388 0.610 0.543 0.192 0.143 0.259 0.205 0.343 0.279 0.290 0.197 0.302 0.214 0.321 0.223
QUAD 0.329 0.266 0.479 0.416 0.646 0.572 0.204 0.155 0.282 0.228 0.385 0.318 0.307 0.210 0.334 0.241 0.379 0.277
Transformer-based without GNNs Methods
HyTrans 0.313 0.255 0.458 0.397 0.621 0.557 0.192 0.140 0.268 0.215 0.372 0.316 0.298 0.204 0.325 0.234 0.361 0.266
GRAN 0.322 0.269 0.472 0.419 0.647 0.593 0.201 0.156 0.287 0.244 0.403 0.349 0.307 0.212 0.326 0.237 0.382 0.290
HyNT 0.295 0.238 0.438 0.377 0.595 0.525 0.175 0.130 0.243 0.192 0.329 0.269 0.285 0.195 0.298 0.216 0.352 0.261
HAHE 0.301 0.239 0.432 0.368 0.607 0.537 0.188 0.142 0.263 0.213 0.367 0.304 0.311 0.221 0.330 0.248 0.387 0.293
HyperFormer 0.338 0.280 0.492 0.434 0.666 0.611 0.213 0.161 0.298 0.255 0.426 0.373 0.352 0.254 0.411 0.325 0.478 0.396
TABLE IV: Evaluation of different models under the fixed-percentage fixed-qualifier scenario on the WikiPeople and JF17K datasets. Best scores are highlighted in bold, the second best scores are underlined. QUAD denotes the QUAD (Parallel) model in Table II, and GRAN denotes the GRAN-hete model in Table II.
Methods WikiPeople-3 WikiPeople-4 JF17K-3 JF17K-4
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
Embedding-based Methods
ShrinkE 0.366 0.291 0.397 0.515 0.222 0.155 0.252 0.353 0.712 0.648 0.745 0.837 0.779 0.726 0.809 0.886
Transformer-based with GNNs Methods
StarE 0.401 0.310 0.434 0.592 0.243 0.156 0.269 0.430 0.707 0.635 0.744 0.847 0.723 0.669 0.753 0.839
QUAD 0.403 0.321 0.438 0.563 0.251 0.167 0.280 0.425 0.730 0.660 0.767 0.870 0.787 0.730 0.823 0.895
Transformer-based without GNNs Methods
HyTrans 0.403 0.323 0.436 0.569 0.248 0.165 0.275 0.422 0.690 0.617 0.725 0.837 0.773 0.717 0.806 0.875
GRAN 0.397 0.328 0.429 0.533 0.239 0.178 0.261 0.364 0.779 0.724 0.811 0.893 0.798 0.744 0.830 0.904
HyNT 0.383 0.306 0.415 0.533 0.233 0.166 0.257 0.369 0.733 0.664 0.771 0.873 0.779 0.725 0.809 0.881
HAHE 0.397 0.323 0.432 0.537 0.251 0.179 0.286 0.392 0.762 0.701 0.791 0.892 0.736 0.666 0.773 0.869
HyperFormer 0.573 0.511 0.603 0.693 0.393 0.336 0.415 0.496 0.832 0.790 0.855 0.914 0.857 0.811 0.884 0.937
Fixed-percentage Mixed-qualifier. Table III presents the results on WD50K, WikiPeople, and JF17K under different fixed ratios of relational facts with qualifiers. For each dataset, we construct three subsets (as describe in § 5.1.1, Point 2) containing approximately 33%, 66%, and 100% of facts with qualifiers. We observe that under different proportions of hyper-relational knowledge HyperMono also achieves optimal performance. Particularly, for the WikiPeople dataset, in which HyperMono performs poorly in the mixed-percentage mixed-qualifier scenario, adding a certain percentage of qualifiers information to triples results in HyperMono outperforming all baselines. For example, for the Hits@1 metric, HyperMono respectively achieves improvements over the best performing baseline, HyperFormer, of 1.2% / 0.9% / 0.6% in the 33% / 66% / 100% variants. This confirms that the proportion of hyper-relational knowledge is a major factor influencing the performance over WikiPeople under the mixed-percentage mixed-qualifier condition. In contrast, on the WD50K dataset, no significant additional improvement was obtained. The most likely reason is the insufficient richness of hyper-relational knowledge in WD50K. For example, in WD50k_100 65.23% of the triples only contain one hyper-relation pair, which limits the functionality of qualifier monotonicity module.
Fixed-percentage Fixed-qualifier. To better understand the impact of hyper-knowledge on the performance, we consider the WikiPeople and JF17K datasets with a fixed number of qualifiers, cf. Table IV. We can observe that HyperMono consistently achieves optimal performance on almost all datasets. Specifically, in terms of the MRR metric, HyperMono respectively shows improvements of 3.5% and 2.4% on the JF17K-3 and JF17K-4 subsets when compared with HyperFormer. Furthermore, we find that for all models, their performance in the fixed-percentage fixed-qualifier setting is generally higher than in the fixed-percentage mixed-qualifier one. The main reason is that in the fixed-percentage fixed-qualifier scenario, the number of qualifier pairs for each triple is fixed, and there is no imbalance in the distribution of hyper-relational knowledge, which further makes the model training process more stable. Therefore, we can conclude that, in general, ensuring a relative stability of hyper-relational knowledge for each triple positively impacts the performance.
5.3 Ablation Studies
To address RQ2, we conduct ablation experiments on WikiPeople-3, WikiPeople-4, JF17K-3, and JF17K-4 subsets to verify the contribution of each component of HyperMono. These include the following three aspects: a) whether the knowledge of qualifiers is considered to infer missing entities, see rows “w/o CNA+TP” and “w/o FNA+QMP” in Table V; b) different neighbor aggregation strategies, see rows “w/o LEI” and “w/o GEI” in Table V; c) without the cone shrink operation in QMP, see row “w/o CSB” in Table V.
TABLE V: Ablation studies about different components. Best scores are highlighted in bold.
Methods WikiPeople-3 WikiPeople-4 JF17K-3 JF17K-4
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
w/o CNA+TP 0.559 0.493 0.591 0.686 0.367 0.305 0.387 0.484 0.832 0.792 0.855 0.912 0.860 0.820 0.885 0.940
w/o GEI 0.584 0.532 0.608 0.690 0.396 0.355 0.404 0.471 0.866 0.838 0.878 0.926 0.883 0.852 0.900 0.949
TABLE VI: Evaluation of the knowledge hypergraph completion task in the mixed arity setting on the N-WikiPeople, N-JF17K and N-FB-AUTO datasets. Best scores are highlighted in bold, the second best scores are underlined.
Methods N-WikiPeople N-JF17K N-FB-AUTO
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
RAE [14] 0.253 0.118 0.343 0.463 0.396 0.312 0.433 0.561 0.703 0.614 0.764 0.854
NaLP [41] 0.338 0.272 0.364 0.466 0.310 0.239 0.334 0.450 0.672 0.611 0.712 0.774
HINGE [15] 0.333 0.259 0.361 0.477 0.473 0.397 0.490 0.618 0.678 0.630 0.706 0.765
NeuInfer [31] 0.350 0.282 0.381 0.467 0.451 0.373 0.484 0.604 0.737 0.700 0.755 0.805
HypE [26] 0.263 0.127 0.355 0.486 0.494 0.399 0.532 0.650 0.804 0.774 0.824 0.856
RAM [30] 0.380 0.279 0.455 0.539 0.539 0.463 0.573 0.690 0.830 0.803 0.851 0.876
HyperMLN [45] 0.351 0.270 0.394 0.497 0.556 0.482 0.597 0.717 0.831 0.803 0.851 0.877
tNaLP+ [46] 0.339 0.269 0.369 0.473 0.449 0.370 0.484 0.598 0.729 0.645 0.748 0.826
S2S [47] 0.372 0.277 0.439 0.533 0.528 0.457 0.570 0.690 - - - -
HyConvE [48] 0.362 0.275 0.388 0.501 0.590 0.478 0.610 0.729 0.847 0.820 0.872 0.901
MSeaKG [32] 0.392 0.290 0.468 0.553 0.561 0.475 0.591 0.705 - - - -
HJE [49] 0.450 0.375 0.487 0.582 0.590 0.507 0.613 0.729 0.872 0.848 0.886 0.903
PolygonE [50] 0.431 0.334 0.454 0.568 0.565 0.485 0.602 0.708 0.858 0.826 0.871 0.921
HyCubE [51] 0.438 0.358 0.482 0.575 0.581 0.507 0.615 0.731 0.884 0.863 0.897 0.920
Different Variants of HyperMono. In Table V, we observe that under most metrics removing any functional module will bring some performance degradation, which justifies the necessity of each component. Specifically:
  1. 1.
    removing either CNA+TP, which are related to coarse-grained answer prediction, or FNA+QMP, which are related to with fine-grained answer prediction, will result in a significant loss of accuracy. The intuitive reason is that the coarse-grained answer set can be used to constrain the decision space of fine-grained answers, preventing the occurrence of answer drift. Meanwhile, the fine-grained answer set allows to refine the answer set, ensuring the reliability of the predicted results.
  2. 2.
    The neighborhood information of known entities has a positive effect on predicting missing entities, and the use of LEI for integrating the neighbors of known entities has a greater impact on inference than GEI.
  3. 3.
    Removing the CSB module will bring a significant decrease in performance. The main reason is that the CSB module captures the qualifier monotonicity of hyper-relational KGs. The main characteristic of the HKGC task is the application of qualifier knowledge in triples to infer answer sets. Through the natural scaling transformation of the cone space, it can be captured that attaching qualifiers to a triple can only narrow down the answer set, but never enlarge it.
5.4 Knowledge Hypergraph Completion
To address RQ3, we further validate the transfer ability of HyperMono. To this aim, we also conduct experiments on the Knowledge Hypergraph Completion (KHC) task. We start by introducing the task definition, dataset descriptions, and baselines. We only introduce necessary background, for a detailed discussion, interested readers are directed to the relevant literatures [47, 48, 32, 49, 50, 51].
Datasets. We test the performance of HyperMono on the KHC task. We conduct experiments on hypergraphs with mixed arity: N-WikiPeople, N-JF17K, and N-FB-AUTO and with fixed arity: N-WikiPeople-3, N-WikiPeople-4, N-JF17K-3, and N-JF17K-4. The statistics of the corresponding datasets are consistent with existing literatures [47, 48, 32, 49, 50, 51].
Baselines. For knowledge hypergraph completion task, we select 14 state-of-the-art baselines, including RAE [14], NaLP [41], HINGE [15], NeuInfer [31], HypE [26], RAM [30], HyperMLN [45], tNaLP+ [46], S2S [47], HyConvE [48], MSeaKG [32], HJE [49], PolygonE [50], and HyCubE [51].
TABLE VII: Evaluation of the knowledge hypergraph completion task in the fixed arity setting on the N-WikiPeople-3, N-WikiPeople-4, N-JF17K-3 and N-JF17K-4 datasets. Best scores are highlighted in bold, the second best scores are underlined.
Methods N-WikiPeople-3 N-WikiPeople-4 N-JF17K-3 N-JF17K-4
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
RAE [14] 0.239 0.168 0.252 0.379 0.150 0.080 0.149 0.273 0.505 0.430 0.532 0.644 0.707 0.636 0.751 0.835
NaLP [41] 0.301 0.226 0.327 0.445 0.342 0.237 0.400 0.540 0.515 0.431 0.552 0.679 0.719 0.673 0.742 0.805
HINGE [15] 0.338 0.255 0.360 0.508 0.352 0.241 0.419 0.557 0.587 0.509 0.621 0.738 0.745 0.700 0.775 0.842
NeuInfer [31] 0.355 0.262 0.388 0.521 0.361 0.255 0.424 0.566 0.622 0.533 0.658 0.770 0.765 0.722 0.808 0.871
HypE [26] 0.266 0.183 - 0.443 0.304 0.191 - 0.527 0.364 0.255 - 0.573 0.408 0.300 - 0.627
RAM [30] 0.254 0.190 - 0.383 0.226 0.161 - 0.367 0.578 0.505 - 0.722 0.743 0.701 - 0.845
HyperMLN [45] 0.252 0.193 - 0.385 0.224 0.167 - 0.370 0.574 0.501 - 0.711 0.734 0.687 - 0.831
tNaLP+ [46] 0.270 0.185 - 0.444 0.344 0.223 - 0.578 0.411 0.325 - 0.617 0.630 0.531 - 0.722
S2S [47] 0.386 0.299 0.421 0.559 0.391 0.270 0.470 0.600 0.740 0.676 0.770 0.860 0.822 0.761 0.853 0.924
HyConvE [48] 0.318 0.240 - 0.482 0.386 0.271 - 0.607 0.729 0.670 - 0.861 0.827 0.770 - 0.931
MSeaKG [32] 0.403 - - 0.579 0.409 - - 0.624 0.754 - - 0.889 0.833 - - 0.938
HJE [49] 0.381 0.303 - 0.596 0.388 0.276 - 0.607 0.707 0.637 - 0.851 0.827 0.775 - 0.928
HyCubE [51] 0.322 0.247 - 0.469 0.370 0.252 - 0.583 0.615 0.544 - 0.752 0.792 0.744 - 0.879
Performance Analysis. The results of the knowledge hypergraph completion experiments in the mixed arity and fixed arity settings are respectively shown in Table VI and Table VII. We observe that our HyperMono consistently achieves state-of-the-art performance on all benchmark datasets. For the mixed arity scenario, compared with the best-performing baseline HyCubE, HyperMono obtains improvements of 4.3% on N-WikiPeople and 9.5% on N-JF17K in the MRR metric. On the N-FB-AUTO dataset, HyperMono virtually matches the results of the best performing baseline. These results demonstrate the capability and effectiveness of HyperMono for capturing and extracting various types of information and knowledge. The same phenomenon occurs in the fixed arity scenario, HyperMono respectively obtains performance improvements of 13.2%, 11.6%, 17.8%, and 9.6% in the MRR metric on the N-WikiPeople-3, N-WikiPeople-4, N-JF17K-3, and N-JF17K-4 datasets, when compared with the best baseline HyCubE. Such a notable performance improvement is mainly due to our explicit treatment of qualifier monotonicity during the encoding process. This feature is more suitable for scenarios where arity is fixed, because introducing qualifier monotonicity can greatly reduce the scope of the optional answer space to obtain the final answer.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 5.4 Knowledge Hypergraph Completion
Figure 4: Parameter sensitivity experiments under different conditions on Wikipeople-3, Wikipeople-4, JF17K-3 and JF17K-4 datasets.
5.5 Parameter Sensitivity

### v2

@@ B   Baselines
  • m-TransH (Wen et al., 2016) builds on TransH by transforming each hyper-relational fact using a star-to-clique conversion.
  • RAE (Zhang et al., 2018) integrates m-TransH with a multi-layer perceptron by considering the relatedness of entities. However, since the distance-based scoring function of TransH enforces constraints on relations, it fails to represent some binary relations in KGs.
  • NaLP-Fix (Rosso et al., 2020) uses a convolutional-based framework to compute a relatedness vector for a triple. However, the model does not distinguish between the main triple and relation-specific qualifiers.
  • HINGE (Rosso et al., 2020) adopts a convolutional framework for modeling hyper-relational facts. However, it only operates on a triple-quintuple level that lacks the necessary granularity for representing a relation instance with its qualifiers.
  • ShrinkE (Xiong et al., 2023) models the primary triplets as a spatial-functional transformation from the head into a relation-specific box.
@@
  • StarE (Galkin et al., 2020) employs a message passing based graph encoder to obtain entity/relation embedings, and feeds these embedings to a transformer decoder to score hyper-relational facts.
  • QUAD (Shomer et al., 2022) introduces a graph encoder that aggregates information from the perspective of the qualifier entities.
@@
  • Hy-Transformer (Yu and Yang, 2021) replaces the graph encoder with light-weight embedding modules, achieving higher efficiency without sacrificing effectiveness.
@@
  • GRAN (Wang et al., 2021) represents a fact as a heterogeneous graph, and employs edge-specific attentive bias to capture both local and global dependencies within the given fact.
@@
  • HyNT (Chung et al., 2023) proposes a unified framework that learns representations of a hyper-relational knowledge graph containing numeric literals in either triples or qualifiers.
@@
  • HAHE (Luo et al., 2023) presents a model with hierarchical attention in global and local level to encode the graphical structure and sequential structure via self-attention layers.
@@
  • HyperFormer (Hu et al., 2023) considers local-level sequential information to encode the content of entities, relations and qualifiers. It includes an entity neighbor aggregator, relation qualifier aggregator and convolution-based bidirectional interaction modules.
@@
Figure 5. Parameter sensitivity experiments under different conditions on Wikipeople-3, Wikipeople-4, JF17K-3 and JF17K-4 datasets.
C   Evaluation Protocol and Implements Details
Evaluation Protocol. Following existing work (Chung et al., 2023; Xiong et al., 2023; Luo et al., 2023; Hu et al., 2023), we rank all entities to predict the head entity in query (?$,r,t,\mathcal{Q})$ or the tail entity in query $(h,r,\texttt{?},\mathcal{Q})$. Specifically, we employ MRR and Hits@$k$ (abbreviated sometimes as H@k, $k\in\{1,3,10\}$) as evaluation metrics. MRR defines the inverse of the rank for the first correct answer, while Hits@k calculates the percentage of correct types ranked among the top-k answers. For all metrics, the larger the value, the better. Following the standard evaluation protocol in the filtered setting (Bordes et al., 2013), all true triples in the HKGs are filtered out during evaluation, since predicting a low rank for these triples should not be penalized. All metrics are computed by averageing over two directions: head entity prediction and tail entity prediction.
Implementation Details. We conduct all experiments on six 32G Tesla V100 GPUs, and employ the AdamW (Loshchilov and Hutter, 2019) optimizer and a cosine decay scheduler with a linear warm-up for optimization. The hyperparameters are tuned according to the MRR on the validation set. We use grid search to select the optimal hyperparameters, mainly including: the learning rate $lr\in$3e-4, 4e-4, 5e-4, 6e-4, 7e-4}, the label smooth values $ls\in$0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9}, the embedding dimension $d\in$50, 100, 150, 200, 400, 600}, the temperature of Equation (1) $\gamma\in$1, 2, 3, 4, 5, 6}, the initialization range of embeddings $ir\in$0.01, 0.02, 0.05, 0.1, 0.2, 0.35, 0.5}, the neighbors of the head entity {2, 3, 4}, the number of qualifiers for a relation {5, 6, 7}, the number of layers in the transformer {2, 4, 8}, the number of heads in transformer {1, 2, 4, 8}, the input dropout rate in the transformer {0.6, 0.7, 0.8}.
D   Parameter Sensitivity